# Unveiling Inefficiencies in LLM-Generated Code: Toward a Comprehensive Taxonomy

### v1

@@ Abstract
Large Language Models (LLMs) are widely adopted for automated code generation with promising results. Although prior research has assessed LLM-generated code and identified various quality issues- such as redundancy, poor maintainability, and sub-optimal performance- a systematic understanding and categorization of these inefficiencies remain unexplored. Without such knowledge, practitioners struggle to optimize LLM-generated code for real-world applications, limiting its adoption. This study can also guide improving code LLMs, enhancing the quality and efficiency of code generation. Therefore, in this study, we empirically investigate inefficiencies in LLM-generated code by state-of-the-art models, i.e., CodeLlama, DeepSeek-Coder, and CodeGemma. To do so, we analyze 492 generated code snippets in the HumanEval++ dataset. We then construct a taxonomy of inefficiencies in LLM-generated code that includes 5 categories (General Logic, Performance, Readability, Maintainability, and Errors) and 19 subcategories of inefficiencies. We then validate the proposed taxonomy through an online survey with 58 LLM practitioners and researchers. Our study indicates that logic and performance-related inefficiencies are the most popular, relevant, and frequently co-occur and impact overall code quality inefficiency. Our taxonomy provides a structured basis for evaluating the quality LLM-generated code and guiding future research to improve code generation efficiency.
@@ I Introduction
While LLMs show promising results in code generation, their generated code, similar to human-written one, exhibit various code quality issues [9, 8], limiting their seamless integration into production environments. For instance, LLM-generated code have shown to be suffered from redundancies, unnecessary computations, and suboptimal implementations [10], resulting in increased execution time, higher memory consumption, maintainability challenges, and import errors [11, 12]. Previous studies have also shown that LLMs may generate incorrect logic, misinterpret task requirements, and struggle to handle corner cases, leading to reliability concerns [13]. Even when functionally correct, LLM-generated code can face quality issues [14], being difficult to maintain [11], or unnecessarily complex [10], prone to security vulnerability and requiring manual intervention before deployment. These issues pose significant risks to software quality, particularly for developers adopting LLMs in real-world settings [15].
While prior research has examined specific quality aspects such as correctness, security, and maintainability in LLM-generated code [16, 17, 14], inefficiencies as a broader concept remain largely unexplored. Existing studies focus only on buggy-generated code [13, 14] to categorize bugs and identify root causes. However, this narrow focus overlooks a critical issue: even functionally correct code can exhibit inefficiencies that degrade software quality, such as performance bottlenecks [10], limiting real-world adoption of LLM-generated code.
In this study, we aim to bridge this gap by identifying, categorizing, and analyzing overall inefficiency patterns, regardless of correctness, in LLM-generated code from three leading open-source LLMs for code generation. We define inefficiencies as any issues in LLM-generated code that degrade its quality, hinder integration into larger systems, or even affect functionality in isolation. Understanding inefficiencies guides the development of more efficient LLMs and helps practitioners and researchers to improve LLM-generated code. Our investigation is done under the following Research Questions (RQs):
@@
We conducted our study using HumanEval+  a widely used benchmark for code generation composed of 164 human-crafted Python tasks [6]. We selected three leading open-source LLM models -CodeLlama, DeepSeek-Coder, and CodeGemma- to generate code for the different tasks, resulting in 492 generated code snippets. To assess LLM-generated code quality and requirements conformity, we employed a higher-capacity GPT-4o-mini model as a judge. The judgments provided initial guidelines for our manual exploratory and qualitative analysis to identify inefficiency patterns in LLM-generated code. Our study results in a taxonomy spanning 5 categories General Logic, Performance, Readability, Maintainability, and Errors, comprising 19 subcategories that is validated by surveying 58 software practitioners and researchers using LLM for coding assistance.
Our findings indicate that General Logic and Performance related inefficiencies are the most frequent, relevant and often co-occur with Errors, Maintainability and Readability and impact the overall efficiency of the code, suggesting that improving LLM reasoning and optimization capabilities could enhance overall code quality. Our survey confirmed these findings, with participants prioritizing correctness, but also expressing concerns about inefficiencies that increase debugging and refactoring efforts.
For LLM developers, our taxonomy offers insights into existing inefficiencies which can inform future targeted improvement directions. For researchers, our categorization provides a structured foundation for further studies on inefficiencies in LLM-generated code. Additionally, our findings can support the development of automated inefficiency detection and refactoring tools, which can enhance code quality as LLM-generated code becomes increasingly adopted in production systems. To summarize, this paper makes the following contributions:
@@
  • We make our data and results publicly available in our replication package. [^1] https://anonymous.4open.science/r/LLMs_Code_Inefficiencies-EEB8
The rest of the paper is organized as follows: In Section II, we present the background. Next, we detail the methodology used to construct and validate the taxonomy of inefficiencies in Section III. In Section IV, we report our empirical findings presenting our taxonomy. Then, we discuss related work and threats in Section V and VI, respectively. Finally, Section VII concludes the paper.
@@ II Background
In this section, we introduce the used three LLMs for code generation in our study, and describe the open coding methodology followed to construct the taxonomy of inefficiencies.
II-A LLMs for Code Generation.
Various LLMs have been introduced for automatic code generation; in our study, we focus on three models as described as follows. CodeLlama is an open-source family of models based on Llama 2 [18] from Meta, designed to support large input contexts and zero-shot instruction following, supporting coding tasks including code completion and infilling on surrounding content. The CodeLlama family includes the foundation model (CodeLlama), a Python-specialized variant (CodeLlama-Python), and an instruction-following model (CodeLlama-Instruct). Each of these models is available in different sizes (7B, 13B, 34B, and 70B parameters). This LLM is primarily trained on a near-deduplicated dataset of publicly available code, with 8% sourced from natural language datasets related to code, including discussions, questions, and answers, while a small portion is sampled to retain natural language understanding skills [4]. In our study, we focused on code completion and utilized the 7B parameter CodeLlama foundation model.
DeepSeek-Coder owns a range of open-source models trained from scratch on 2 trillion tokens across 87 programming languages. The pretraining data is organized at the repository level to enhance model capability. The models employ two key techniques: fill-in-the-middle and next-token prediction. Just like CodeLlama, DeepSeek-Coder is also available in different sizes, ranging from 1.3B to 33B parameters, with both base and instruction-tuned versions. These models leverage a 16K context window to improve code generation and infilling [19]. In our study, we focused on code completion and utilized the 6.7B parameter DeepSeek-Coder base model.
CodeGemma is a collection of specialized open-code models built on top of Google DeepMind’s Gemma models [20], further trained on 500 to 1000 billion tokens, primarily for code-related tasks such as code completion and generation. The CodeGemma family includes three models: a 7B pre-trained model, a 7B instruction-tuned model, and a specialized 2B model designed for code infilling and open-ended generation. The 2B model is trained entirely on code, while the 7B models are trained on a mix of 80% code and 20% natural language, sourced from deduplicated publicly available code repositories. These models are trained using a fill-in-the-middle task to enhance their coding capabilities [21]. In our study, we used the CodeGemma 7B pre-trained model for code completion.
II-B Open Coding.
Open coding is a qualitative research method used to analyze textual data, for example, survey responses and codes, by breaking it down into concepts and categories in an unstructured approach [22]. It consists of three parts: noticing, collecting, and thinking [23]. Initially, researchers examine closely raw data by looking into similarities and dissimilarities, notice patterns, and mark them with appropriate code, i.e. label to define categories based on noticed characteristics and properties. This approach is not linear as researchers often need to revisit data and adjust labels -adding, removing, or merging- to develop a final codebook. Open coding is known to be postformed code since codes are formed incrementally during the process [22]. Postformed codes have proven to be more useful in SE research than preformed labels [24, 25]. During open coding, independent reviewers begin by coding the same set of samples, identifying patterns and categories relevant to the study’s focus. They then engage in discussions to consolidate their findings into an initial codebook, resolving disagreements through communication or by involving a third party. To ensure reliability, the level of agreement between reviewers is measured before discussion [22]. Previous studies in software engineering have employed open coding for qualitative research, including the construction of taxonomies as we do here [26, 24, 25].
@@ III Methodology
This section outlines the methodology we used in this paper to construct and validate the taxonomy of inefficiencies in LLM-generated code, consisting of 5 steps, as illustrated in Figure 1. First, we collected data by generating code using the HumanEval+ benchmark [6] with three open-source LLM4Code models: CodeLlama [4], DeepSeek-Coder [19], and GodeGemma [21]. Second, we leveraged a higher-capacity model (GPT-4o-mini) to judge the generated code and identify existing inefficiencies. In the third and fourth steps, we manually analyzed the code samples and the model’s judgments to develop a comprehensive taxonomy. Finally, in the fifth step, we surveyed practitioners to validate the LLM-generated code inefficiencies in our taxonomy. We describe details of each step in the rest of this section.
@@ III-A Step 1: Data Selection and Collection
Regarding the data collection, considering the popularity and wide adoption of Python as a programming language[27], this study focused on Python-generated LLM code. Given the focus on Python, we selected HumanEval+ [6] (denoted as $\mathit{A}$ in Figure 1), a widely recognized and extensively employed benchmark for code generation [28, 29, 30, 31]. HumanEval+ enhances the original HumanEval [32]dataset by augmenting it with additional test cases, enabling a more rigorous and reliable assessment of functional correctness [6]. Such a benchmark comprises 146 Python programming tasks, each accompanied by (i) a task description provided as a doc string and (ii) associated unit tests to rigorously evaluate the correctness of the code snippets. Regarding the LLMs, we aimed to select those trained or fine-tuned for code generation tasks to enable a more meaningful and targeted analysis of inefficiencies. To ensure accessibility and feasibility, we established two selection criteria: (i) the models must be open source and (ii) have fewer than 8B parameters. Based on these criteria, we selected the three highest-performing models according to the HumanEval+ leaderboard [33] [^2] Accessed July 10, 2024: CodeLlama-7B [^3] https://huggingface.co/meta-llama/CodeLlama-7b-hf, GodeGemma-7B [^4] https://huggingface.co/google/codegemma-7b, and DeepSeek Coder-6.7B [^5] https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-base.
Once defined the setup for our study, we moved forward by generating the code snippets. Regarding the hyperparameters, we adopted a conservative approach, setting the temperature to 00, prioritizing deterministic outputs, and reducing randomness in code generation. For prompt construction, we adhered to default configuration from official Hugging Face documentation. This approach ensured that the models could generate code efficiently and accurately within the defined constraints. As a result of this step, we have a dataset of $164$ generated code snippets for each LLM, resulting in $492$ code snippets ($\mathit{B}$), from the initial dataset of tasks ($\mathit{A}$).
@@ III-B Step 2: Judging LLM-generated Code
LLMs have been effectively used to assess code or tests [34, 35], proposing a paradigm where LLMs serve as judges (LaaJ). Prior studies have also leveraged LLM to assist in identifying a generated-code bug’s root causes [14]. In this context, we propose leveraging LLMs to assess the quality of generated code. Specifically, we employed the GPT-4o-mini model, which offers strong performance comparable to GPT-4 while being more cost-effective [36]. Regarding the hyperparameters, we considered a temperature of $\mathit{0}$ aiming to achieve a more conservative generation and max_tokens set to $\mathit{5120}$ to ensure sufficient space for detailed output. Knowing that GPT-4 models can be shaped and optimized by setting system prompts, which define the model’s persona [37], we adopted a persona of an expert in code quality analysis, with a focus on identifying specific inefficiencies in machine-generated code. Such a persona guides the model prioritizes practical and significant code inefficiencies rather than minor or irrelevant issues. Finally, we defined a fixed user prompt used as input when asking for judgment support (judger). In this prompt, we input the generated code snippet and its task description, providing a context to aid the judger in identifying logical inefficiencies in the generated code. As a result of this step, we enriched the dataset $\mathit{B}$ with GPT-generated judgments, creating dataset $\mathit{C}$, as denoted in Figure 1. Specifically, $\mathit{C}$ consists of 492 generated code samples, each paired with its corresponding LLM-generated judgment.
@@ III-C Step 3 and 4: Manual Labeling
Step 3: Manual Labeling - Pilot Study With no predefined categories in previous studies, our objective is to construct a taxonomy of inefficiencies using an inductive, bottom-up approach through manual analysis of LLM-code generated snippets. To initiate this process, we randomly selected 196 samples $\mathit{C}$ (40%, pilot set), ensuring an equal proportion of cases from each model. The first two authors, senior Master’s student and a Postdoctoral Fellow with experience in Python and SE research, collaboratively analyzed and labeled the pilot set (40% of samples) using open coding [22]. The analysis involved examining the task description, the generated code, code quality judgment, and the execution and testing status (from the unit tests available in HumanEval+). Without predefined assumptions, inefficiencies were iteratively identified and grouped into categories and subcategories. Knowing that LLMs as judgers may hallucinate [35], we used their judgments as a starting point but independently verified and expanded the inefficiencies. During this process, the labelers assigned brief, descriptive labels to each inefficiency observed in the pilot set. These initial labels were then categorized, forming a hierarchical taxonomy of inefficiency types. This iterative process involved continuously refining the categories as they moved back and forth between individual samples and emerging patterns, ensuring that the taxonomy accurately reflected the inefficiencies underlying generated code. All tasks in the pilot set were thoroughly discussed, and the labelers reached a consensus on both categories and subcategories, resulting in a taxonomy comprising $5$ inefficiency categories and $18$ subcategories and a fully labeled pilot set. This step resulted a codebook of inefficiency patterns, with multi labeling allowed as a single code snippet can exhibit multiple inefficiencies [13].
@@
Step 4: Taxonomy Refinement & Full Dataset Labeling After labeling the pilot set and constructing the initial codebook of inefficiencies, the labelers independently labeled the remaining 60% samples, following the same procedure. If any generated code could not be categorized within the existing taxonomy, they were marked Pending and reviewed to determine if new categories or subcategories should be created to accommodate these cases. Discussions, led to adding Syntax Error as a subcategory in the Errors category, resulting in a final taxonomy of 5 categories and 19 subcategories to better capture inefficiencies observed in LLM-generated code. Once all samples were labeled, we assessed inter-rater agreement to evaluate the consistency between the two labelers. Following similar studies [26, 13], we used Cohen’s Kappa [38], calculated at category and subcategory levels. Inter-rater agreement (Cohen’s Kappa) reached 0.846 (category) and 0.735 (subcategory), aligning with prior SE studies[26, 13]. Posts labeled as ‘Pending‘ were included in the computation of Cohen’s Kappa only when both raters agreed on this designation. Labeling disagreements were resolved through discussion, leading to a consensus on the final labeling.
III-D Step 5: Participants feedback and survey analysis

### v2

@@ Abstract
Large Language Models (LLMs) are widely adopted for automated code generation with promising results. Although prior research has assessed LLM-generated code and identified various quality issues- such as redundancy, poor maintainability, and sub-optimal performance- a systematic understanding and categorization of these inefficiencies remain unexplored. Therefore, we empirically investigate inefficiencies in LLM-generated code by state-of-the-art models, i.e., CodeLlama, DeepSeek-Coder, and CodeGemma. To do so, we manually analyze 492 generated code snippets in the HumanEval+ dataset. We then construct a taxonomy of inefficiencies in LLM-generated code that includes 5 categories (General Logic, Performance, Readability, Maintainability, and Errors) and 19 subcategories of inefficiencies. We validate the obtained taxonomy through an online survey with 58 LLM practitioners and researchers. The surveyed participants affirmed the completeness of the proposed taxonomy, and the relevance and the popularity of the identified code inefficiency patterns. Our qualitative findings indicate that inefficiencies are diverse and interconnected, affecting multiple aspects of code quality, with logic and performance-related inefficiencies being the most frequent and often co-occur while impacting overall code quality. Our taxonomy provides a structured basis for evaluating the quality of LLM-generated code and guiding future research to improve code generation efficiency.
@@ I Introduction
While LLMs show promising results in code generation, their generated code, similar to human-written code, exhibit various quality issues [9, 8], which can impede their smooth integration into production environments. For instance, LLM-generated code has shown to suffer from redundancies, unnecessary computations, and suboptimal implementations [10], resulting in increased execution time, higher memory consumption, maintainability challenges, and import errors [11, 12]. Previous studies have also shown that LLMs may generate incorrect logic, misinterpret task requirements, and struggle to handle corner cases, leading to reliability issues [13]. Even when functionally correct, LLM-generated code can face quality issues [14]. It may be difficult to maintain [11], unnecessarily complex [10], prone to security vulnerabilities, and require manual intervention before deployment. These issues pose significant risks to software quality, particularly for developers adopting LLMs in real-world settings [15].
While prior research has examined specific quality aspects, such as correctness, security, and maintainability in LLM-generated code [16, 17, 14], inefficiencies as a broader concept remain largely unexplored. Existing studies focus only on buggy-generated code [13, 14] to categorize bugs and identify root causes. However, this narrow focus overlooks a critical issue: even functionally correct code can exhibit inefficiencies that degrade software quality, such as performance bottlenecks [10], limiting real-world adoption of LLM-generated code.
In this study, we aim to bridge this gap by identifying, categorizing, and analyzing overall inefficiency patterns, regardless of correctness, in LLM-generated code from three leading open-source LLMs for code generation. In this work, we define inefficiencies as any issues in LLM-generated code that degrade its quality, hinder integration into larger systems, or even affect functionality in isolation. Understanding inefficiencies guides the development of more efficient LLMs and helps practitioners and researchers improve LLM-generated code. Our investigation is guided by the following Research Questions (RQs):
@@
We conduct our study using HumanEval+  a widely used benchmark for code generation composed of 164 human-crafted Python tasks [6]. We selected three leading open-source LLM models -CodeLlama, DeepSeek-Coder, and CodeGemma- to generate code for the different tasks, resulting in 492 generated code snippets. To evaluate the quality and requirements conformity of the LLM-generated code, we employed the higher-capacity GPT-4o-mini model as a judge. The judgments provided initial guidelines for our manual exploratory and qualitative analysis to identify inefficiency patterns in LLM-generated code. We adopted open coding [18] to iteratively derive categories of inefficiencies. Our study results in a taxonomy spanning five categories: General Logic, Performance, Readability, Maintainability, and Errors, comprising 19 subcategories. We validate this taxonomy through a survey with 58 software practitioners and researchers who use LLMs for coding assistance.
Our findings indicate that 33.54% of the studied sample exhibited multiple inefficiencies, indicating that inefficiencies in LLM-generated code are diverse and interconnected. General Logic and Performance inefficiencies are the most frequent and often co-occurring with Maintainability, and Readability inefficiencies. This co-occurrence suggests that limitations in LLMs’ ability to generate correct and efficient code contribute to broader inefficiencies affecting code quality. The surveyed participants largely validated the relevance and prevalence of identified inefficiencies, confirming that it is aligned with real-world inefficiencies in LLM-generated code regardless of the model used. Additionally, while participants acknowledged these inefficiencies, they emphasized that in practice, they prioritize correctness over other code aspects. Our findings highlight a critical gap in LLMs’ capability to generate correct, optimized, and high-quality code.
For LLM developers, our taxonomy offers insights into existing inefficiencies of LLM-generated code. This knowledge can inform future targeted LLMs code generation improvement directions. For researchers, our categorization provides a structured foundation for further studies on inefficiencies in LLM-generated code. Additionally, our findings can support the development of automated detection techniques and refactoring tools targeting these inefficiencies to enhance code quality as LLM-generated code becomes increasingly adopted in production systems. To summarize, this paper makes the following contributions:
@@
  • We make our data and results publicly available in our replication package [19].
The rest of the paper is organized as follows: In Section II, we present the background. Next, we detail the methodology used to construct and validate the taxonomy of inefficiencies in Section III. In Section IV, we report our empirical findings, presenting our taxonomy. Then, we discuss related work and threats to the validity of our results in Section V and VI, respectively. Finally, Section VII concludes the paper.
@@ II Background
In this section, we introduce the three LLMs used for code generation in our study.
CodeLlama
is an open-source family of models based on Llama 2 [20] developed by Meta for code-related tasks such as completion and infilling. It includes the foundation model (CodeLlama), a Python-specialized model (CodeLlama-Python), and an instruction-following model (CodeLlama-Instruct), available in different sizes (7B, 13B, 34B, and 70B parameters). CodeLlama is trained on a near-deduplicated dataset of publicly available code, it incorporates a small fraction of natural language data related to programming to enhance comprehension  [4].
DeepSeek-Coder
provides a range of open-source models trained from scratch on 2 trillion tokens across 87 programming languages. The pretraining data is organized at the repository level to enhance model capability [21]. The models employ two key techniques: fill-in-the-middle and next-token prediction. Just like CodeLlama, DeepSeek-Coder is also available in different sizes, ranging from 1.3B to 33B parameters, with both foundational and instruction-tuned versions. These models leverage a 16K context window to improve code generation and infilling [21].
CodeGemma
is a collection of specialized open-code models based on Google DeepMind’s Gemma models [22], trained on 500 to 1000 billion tokens, primarily for code-related tasks. The CodeGemma family includes: a 7B pre-trained model, a 7B instruction-tuned model, and a specialized 2B model designed for code infilling and open-ended generation. The 2B model is trained entirely on code, while the 7B models are trained on a mix of 80% code and 20% natural language, sourced from deduplicated publicly available code repositories. These models are trained using a fill-in-the-middle task to enhance their coding capabilities [23].
@@ III Methodology
This section presents the methodology used in this study to construct and validate the taxonomy of inefficiencies in LLM-generated code. The process consists of five steps, as illustrated in Figure 1. First, we generated code using the HumanEval+ benchmark [6] and collected data from three open-source LLM4Code models: CodeLlama [4], DeepSeek-Coder [21], and CodeGemma [23]. Second, we leveraged a higher-capacity model (GPT-4o-mini) to judge the generated code and identify existing inefficiencies. In the third and fourth steps, we manually analyzed the code samples and the model’s judgments to develop a comprehensive taxonomy. Finally, in the fifth step, we conducted a survey with practitioners and researchers to validate the LLM-generated code inefficiencies identified in our taxonomy. The details of each step are presented in the remainder of this section.
@@ III-A Step 1: Data Selection and Collection
Regarding the data collection, considering the popularity and wide adoption of Python as a programming language [24], this study focused on Python-generated LLM code. Given the focus on Python, we selected HumanEval+ [6] (denoted as $\mathit{A}$ in Figure 1), a widely recognized and extensively employed benchmark for code generation [25, 26, 27, 28]. HumanEval+ enhances the original HumanEval [29] dataset by augmenting it with additional test cases, enabling a more rigorous and reliable assessment of the functional correctness of the generated code [6]. HumanEval+ comprises 146 Python programming tasks, each accompanied by (i) a task description provided as a doc string and (ii) associated unit tests to rigorously evaluate the correctness of the code snippets. Regarding the LLMs, we aimed to select those trained or fine-tuned for code generation tasks to enable a more meaningful and targeted analysis of inefficiencies. To ensure accessibility and feasibility, we established two selection criteria: (i) the models must be open source and (ii) have fewer than 8B parameters. Based on these criteria, we selected the three highest-performing models according to the HumanEval+ leaderboard [30]: CodeLlama-7B [31], GodeGemma-7B [32], and DeepSeek Coder-6.7B [33]. These models, beyond ranking among the top-performing models [30], are also widely adopted for code generation tasks [27]. Although our selection focuses on small open-source models, these models demonstrate performance comparable to proprietary ones, surpassing GPT-3.5 [21] and approaching GPT-4 [34], ensuring representativeness in our study. After defining the setup for our study, we proceeded with generating the code snippets. For the hyperparameters, we adopted a conservative approach by setting the temperature to 00, ensuring deterministic outputs and minimizing randomness in code generation. For prompt construction, we followed the default configuration recommended in the official Hugging Face documentation [32, 31, 33]. This approach ensured a fair assessment of inherent inefficiencies in LLM-generated code rather than artifacts caused by inadequate prompts. As a result of this step, we have a dataset of $164$ generated code snippets for each LLM, resulting in $492$ code snippets ($\mathit{B}$), from the initial dataset of tasks ($\mathit{A}$).
@@ III-B Step 2: Judging LLM-generated Code
LLMs have been effectively used to assess code or tests [35, 36], proposing a paradigm where LLMs serve as judges (LaaJ). Prior studies have also leveraged LLMs to assist in identifying a generated-code bug’s root causes [14]. In this context, we propose to leverage LLMs to assess the quality of generated code. Specifically, we employ the GPT-4o-mini model, which offers performance comparable to GPT-4 while being more cost-effective [37]. Regarding the hyperparameters, we considered a temperature of $\mathit{0}$ aiming to achieve a more conservative generation and max_tokens set to $\mathit{5120}$ to ensure sufficient space for detailed output.
Knowing that GPT-4 models can be shaped and optimized by setting system prompts, which define the model’s persona [38], we adopted the persona of an expert in code quality analysis with a focus on identifying specific inefficiencies in machine-generated code. Such a persona guides the model to prioritize practical and significant code inefficiencies rather than minor or irrelevant inefficiencies.
Finally, we defined a fixed user prompt used as input when asking for judgment support (judger). In this prompt, we input the generated code snippet and its task description, providing a context to aid the judger in identifying logical inefficiencies in the generated code. As a result of this step, we enriched the dataset $\mathit{B}$ with GPT-generated judgments, creating dataset $\mathit{C}$, as denoted in Figure 1. Specifically, $\mathit{C}$ consists of 492 generated code samples, each paired with its corresponding LLM-generated judgment.
@@ III-C Step 3 and 4: Manual Labeling
Step 3: Manual Labeling - Pilot Study With no predefined categories from previous studies, our objective is to construct a taxonomy of inefficiencies using an inductive, bottom-up approach through manual analysis of LLM-generated code snippets. To initiate this process, we randomly selected 196 samples $\mathit{C}$ (40%, pilot set), ensuring an equal proportion of cases from each model. The first two authors, a senior Master’s student and a Postdoctoral Fellow with experience in Python and Software Engineering (SE) research, collaboratively analyzed and labeled the pilot set (40% of samples) using open coding [18], a method widely used for taxonomy construction in SE [39, 40, 41]. The analysis involved examining the task description, the generated code, code quality judgment, and the execution and testing status (from the unit tests available in HumanEval+). Without predefined assumptions, inefficiencies were iteratively identified and grouped into categories and subcategories. Since LaaJ may hallucinate [36], we treated their judgments as a starting point and independently verified and expanded on the identified inefficiencies. During this process, the labelers assigned brief, descriptive labels to each inefficiency observed in the pilot set. These initial labels were then categorized, forming a hierarchical taxonomy of inefficiency types. This iterative process involved continuously refining the categories as they moved back and forth between individual samples and emerging patterns, ensuring that the taxonomy accurately reflected the inefficiencies underlying the generated code. All tasks in the pilot set were thoroughly discussed, and the labelers reached a consensus on both categories and subcategories, resulting in a taxonomy comprising $5$ inefficiency categories and $18$ subcategories and a fully labeled pilot set. This step resulted in a codebook of inefficiency patterns, with multi-labeling allowed, since a single code snippet can exhibit multiple types of inefficiencies [13].
@@
Step 4: Taxonomy Refinement & Full Dataset Labeling After labeling the pilot set and constructing the initial codebook of inefficiencies, the labelers independently labeled the remaining 60% samples, following the same procedure. If any generated code could not be categorized within the existing taxonomy, they were marked as Pending and reviewed to determine whether new categories or subcategories were needed. Through these discussions, Syntax Error was added as a subcategory under Errors, leading to a refined taxonomy comprising five categories and 19 subcategories, providing a more comprehensive representation of inefficiencies in LLM-generated code.
Once all samples were labeled, we assessed inter-rater agreement to evaluate the consistency between the two labelers. Following similar studies [39, 13], we used Cohen’s Kappa [42], calculated at category and subcategory levels. Inter-rater agreement (Cohen’s Kappa) reached 0.846 (for category) and 0.735 (for subcategory), aligning with prior SE studies [39, 13].
III-D Step 5: Participants’ feedback and survey analysis